# Sensor Informativeness, Identifiability, and Uncertainty in Bayesian Inverse Problems for Structural Health Monitoring

### v1

@@ 2 Forward Problem and Inverse Formulation
Partition $[0,L]$ into NN elements with endpoints $\{s_{j}\}_{j=0}^{N}$, lengths $\Delta s_{j}:=s_{j}-s_{j-1}$, and midpoints $\{\tilde{s}_{j}\}_{j=1}^{N}$. Define the element value of the (unknown) compliance as the element average
@@ 3.1 Likelihood
where $u:=F(\mathbf{m})-\mathbf{y}$ denotes the residual vector. Here $F(\mathbf{m})$ is the discrete forward map. In the linear case $F(\mathbf{m})=A\mathbf{m}$ with design matrix AA, so (14) reduces to the familiar weighted least–squares setting [tarantolaInverseProblemTheory2005]. The matrix $\Gamma\succ 0$ encodes relative correlations between measurements (for instance across load positions at a fixed sensor), while the scalar $\sigma^{2}$ scales that correlation to the overall noise level.
@@ 3.4 Computation for nonlinear forward models
and replace AA by $J_{k}$ in (17) to obtain a local Gaussian posterior. Iterating these updates yields a Gauss–Newton (GN) or Levenberg–Marquardt (LM) sequence that converges to the MAP estimator under standard conditions. At the MAP, a Laplace approximation supplies a covariance
@@ 4 Sensor informativeness via Fisher information
Quantifying the information content of a sensor at location $r$ regarding the flexural rigidity EIEI at position $x$ is crucial for optimizing experimental design. Such an assessment guides sensor placement by indicating where measurements are most informative about the flexural rigidity field. We adopt the Fisher information matrix (FIM) [fisherMathematicalFoundationsTheoretical1922, haraldcramerMathematicalMethodsStatistics1946] as a local, model-based measure of identifiability: it quantifies how strongly the data constrain different directions of the parameter vector. Under the Gaussian likelihood in (14), the (expected) FIM is
@@
Because the measurements are stacked by sensor and load, it is convenient to partition the Jacobian as $J=\big[J_{1}^{\top}\,\cdots\,J_{R}^{\top}\big]^{\top}$ with $J_{i}\in\mathbb{R}^{K\times N}$ collecting the rows associated with sensor ii. With the same ordering, the noise covariance is block diagonal, $\Gamma=\mathrm{diag}(\Gamma_{1},\dots,\Gamma_{R})$, where $\Gamma_{i}\in\mathbb{R}^{K\times K}$ stores the correlations for sensor ii. The information then decomposes additively as:
@@
where $A_{(i,k),j}$ is the element integral of the analytic kernel for sensor $r_{i}$ and load position $z_{k}$ (see (12)). We refer to the map $x_{j}\mapsto\mathcal{I}^{(i)}_{jj}$ as the informativeness curve of sensor ii. Interpreted through the Cramér–Rao bound (and ignoring cross-parameter couplings), larger diagonal values correspond to smaller approximate local variances for the associated flexural rigidity component.
@@ 5.1 Ill-posedness and sources of uncertainty
The recovery of the flexural rigidity from rotation data is an ill-posed inverse problem in the sense of Hadamard: the forward operator AA in (11) is compact and its singular values decay rapidly. Two implications are central for interpretation. First, noise amplification: small perturbations in the measured data $\mathbf{y}$ can induce large changes in the estimated flexural rigidity field. Second, smoothing (low-pass) behaviour: fine-scale features of $EI(x)$ are weakly observable and can only be reconstructed with large uncertainty.
@@ 5.2 Spatial variation of identifiability
On a simply supported span under traversing point loads, the bending moment vanishes at the supports. In the forward relation (4), the integrand $m_{r}(s)M(s;z)v(s)$ therefore contributes little near the ends across all load positions $z$, so the corresponding columns of AA have small norms. As a consequence, the per-element Fisher information (22) is minimal near $x=0$ and $x=L$, and the posterior variance is correspondingly large in those regions. Toward mid-span, sensitivities increase and uncertainty narrows. Between widely separated measurement stations, however, distinct flexural rigidity profiles can produce nearly indistinguishable rotation responses, leading to practical non-identifiability of the flexural rigidity distribution in those intervals. These spatial patterns of information content are consistent with the per-sensor informativeness curves discussed in §4 and are illustrated in Fig.˜4.
@@
Figure 6: Bayesian identification of flexural rigidity $EI(x)$ from rotation (tilt) data using 10 elements. Posterior EIEI profiles for decreasing rotation-noise levels. Black step: ground-truth EIEI (with an artificial damage zone); blue line: posterior mean; shaded bands: $75\%$ (dark) and $95\%$ (light) credible intervals. As $\sigma$ decreases, uncertainty shrinks and the location/magnitude of the flexural rigidity reduction are recovered more sharply; improvements are most pronounced away from the supports, where the moment field affords greater leverage on EIEI.
@@ 5.3 Bias–Variance Trade-off in Flexural-Rigidity Recovery
For any fixed discretization NN, number of stations $R$, and regularization weight $\lambda$ (recall $\lambda=\sigma^{2}\tau$), the total mean–squared error at mid-span ($x=L/2$) decomposes into a variance term and a bias term. Writing $\widehat{EI}_{1/2}=\widehat{EI}(L/2)$ and letting $EI^{\star}(x)$ denote the true flexural rigidity,
@@
Figure 7 shows $\mathrm{RMSE}\penalty 10000\ \!\big(\widehat{EI}_{1/2}\big)$ versus NN for several $R$. Three regimes are visible.
Coarse meshes (small NN): the field is under-resolved and the $D_{2}$ prior enforces strong smoothing. The bias term in (23) dominates, so RMSE is large. Refining the mesh reduces discretization bias and the curves fall.
@@
Over-parameterized meshes (large NN): The number of unknowns grows while the data information saturates ($K$ and $R$ are fixed). The system becomes increasingly ill-conditioned as high-frequency modes become unconstrained by the data. Consequently, noise propagation amplifies the variance term in (23), and the RMSE increases accordingly.
Increasing the number of sensors $R$ raises the data precision $\sum_{i}J_{i}^{\top}\Gamma_{i}^{-1}J_{i}$ (see §4), which lowers the variance across all NN and reduces reliance on strong regularization, thereby also trimming bias. In Fig. 7 the curves shift downward nearly in parallel as $R$ grows, and the minimum RMSE occurs in the same NN band: discretization bias is governed by element size, so added sensors chiefly compress the variance without moving the bias-controlled optimum far to the right.
In Fig. 7, $\lambda$ is selected once per $R$ by the Quasi-Optimality criterion and then reused along the NN–sweep. This mimics practice (a single tuning per instrumentation layout). If $\lambda$ were re-tuned at every $(N,R)$, the right-hand rise would be partially damped (more regularization at larger NN), but the qualitative U-shape set by (23) would remain.
The mid-span error is governed by the balance in (23): refine until discretization bias no longer dominates; beyond that point, without additional information (larger $R$ or richer load sets), the variance necessarily grows and RMSE worsens. This provides a practical criterion for choosing NN given a sensing plan.
@@ 6 Field application: openLAB research bridge
Fig.˜12 shows the reconstruction using six piecewise–constant elements. Each EIEI value therefore represents an element average, and $75\%$ and $95\%$ credible bands quantify uncertainty. The posterior mean indicates lower flexural rigidity in span 1 than in span 2, consistent with construction records. While both spans possess identical cross-sections, Span 1 was cast with concrete strength class C25/30 and Span 2 with C50/60, implying a lower elastic modulus for the former. Crucially, the uncertainty is spatially heterogeneous: the bands remain narrow in the span interiors but widen significantly near the zero-moment zones, reflecting reduced sensor informativeness in those regions.

### v2

@@ 2 Forward Problem and Inverse Formulation
Partition $[0,L]$ into $N$ elements with endpoints $\{s_{j}\}_{j=0}^{N}$, lengths $\Delta s_{j}:=s_{j}-s_{j-1}$, and midpoints $\{\tilde{s}_{j}\}_{j=1}^{N}$. Define the element value of the (unknown) compliance as the element average
@@ 3.1 Likelihood
where $u:=F(\mathbf{m})-\mathbf{y}$ denotes the residual vector. Here $F(\mathbf{m})$ is the discrete forward map. In the linear case $F(\mathbf{m})=A\mathbf{m}$ with design matrix $A$, so (14) reduces to the familiar weighted least–squares setting [tarantolaInverseProblemTheory2005]. The matrix $\Gamma\succ 0$ encodes relative correlations between measurements (for instance across load positions at a fixed sensor), while the scalar $\sigma^{2}$ scales that correlation to the overall noise level.
@@ 3.4 Computation for nonlinear forward models
and replace $A$ by $J_{k}$ in (17) to obtain a local Gaussian posterior. Iterating these updates yields a Gauss–Newton (GN) or Levenberg–Marquardt (LM) sequence that converges to the MAP estimator under standard conditions. At the MAP, a Laplace approximation supplies a covariance
@@ 4 Sensor informativeness via Fisher information
Quantifying the information content of a sensor at location $r$ regarding the flexural rigidity $EI$ at position $x$ is crucial for optimizing experimental design. Such an assessment guides sensor placement by indicating where measurements are most informative about the flexural rigidity field. We adopt the Fisher information matrix (FIM) [fisherMathematicalFoundationsTheoretical1922, haraldcramerMathematicalMethodsStatistics1946] as a local, model-based measure of identifiability: it quantifies how strongly the data constrain different directions of the parameter vector. Under the Gaussian likelihood in (14), the (expected) FIM is
@@
Because the measurements are stacked by sensor and load, it is convenient to partition the Jacobian as $J=\big[J_{1}^{\top}\,\cdots\,J_{R}^{\top}\big]^{\top}$ with $J_{i}\in\mathbb{R}^{K\times N}$ collecting the rows associated with sensor $i$. With the same ordering, the noise covariance is block diagonal, $\Gamma=\mathrm{diag}(\Gamma_{1},\dots,\Gamma_{R})$, where $\Gamma_{i}\in\mathbb{R}^{K\times K}$ stores the correlations for sensor $i$. The information then decomposes additively as:
@@
where $A_{(i,k),j}$ is the element integral of the analytic kernel for sensor $r_{i}$ and load position $z_{k}$ (see (12)). We refer to the map $x_{j}\mapsto\mathcal{I}^{(i)}_{jj}$ as the informativeness curve of sensor $i$. Interpreted through the Cramér–Rao bound (and ignoring cross-parameter couplings), larger diagonal values correspond to smaller approximate local variances for the associated flexural rigidity component.
@@ 5.1 Ill-posedness and sources of uncertainty
The recovery of the flexural rigidity from rotation data is an ill-posed inverse problem in the sense of Hadamard: the forward operator $A$ in (11) is compact and its singular values decay rapidly. Two implications are central for interpretation. First, noise amplification: small perturbations in the measured data $\mathbf{y}$ can induce large changes in the estimated flexural rigidity field. Second, smoothing (low-pass) behaviour: fine-scale features of $EI(x)$ are weakly observable and can only be reconstructed with large uncertainty.
@@ 5.2 Spatial variation of identifiability
On a simply supported span under traversing point loads, the bending moment vanishes at the supports. In the forward relation (4), the integrand $m_{r}(s)M(s;z)v(s)$ therefore contributes little near the ends across all load positions $z$, so the corresponding columns of $A$ have small norms. As a consequence, the per-element Fisher information (22) is minimal near $x=0$ and $x=L$, and the posterior variance is correspondingly large in those regions. Toward mid-span, sensitivities increase and uncertainty narrows. Between widely separated measurement stations, however, distinct flexural rigidity profiles can produce nearly indistinguishable rotation responses, leading to practical non-identifiability of the flexural rigidity distribution in those intervals. These spatial patterns of information content are consistent with the per-sensor informativeness curves discussed in §4 and are illustrated in Fig.˜4.
@@
Figure 6: Bayesian identification of flexural rigidity $EI(x)$ from rotation (tilt) data using 10 elements. Posterior EIEI profiles for decreasing rotation-noise levels. Black step: ground-truth EIEI (with an artificial damage zone); blue line: posterior mean; shaded bands: $75\%$ (dark) and $95\%$ (light) credible intervals. As $\sigma$ decreases, uncertainty shrinks and the location/magnitude of the flexural rigidity reduction are recovered more sharply; improvements are most pronounced away from the supports, where the moment field affords greater leverage on $EI$.
@@ 5.3 Bias–Variance Trade-off in Flexural-Rigidity Recovery
For any fixed discretization $N$, number of stations $R$, and regularization weight $\lambda$ (recall $\lambda=\sigma^{2}\tau$), the total mean–squared error at mid-span ($x=L/2$) decomposes into a variance term and a bias term. Writing $\widehat{EI}_{1/2}=\widehat{EI}(L/2)$ and letting $EI^{\star}(x)$ denote the true flexural rigidity,
@@
Figure 7 shows $\mathrm{RMSE}\penalty 10000\ \!\big(\widehat{EI}_{1/2}\big)$ versus $N$ for several $R$. Three regimes are visible.
Coarse meshes (small $N$): the field is under-resolved and the $D_{2}$ prior enforces strong smoothing. The bias term in (23) dominates, so RMSE is large. Refining the mesh reduces discretization bias and the curves fall.
@@
Over-parameterized meshes (large $N$): The number of unknowns grows while the data information saturates ($K$ and $R$ are fixed). The system becomes increasingly ill-conditioned as high-frequency modes become unconstrained by the data. Consequently, noise propagation amplifies the variance term in (23), and the RMSE increases accordingly.
Increasing the number of sensors $R$ raises the data precision $\sum_{i}J_{i}^{\top}\Gamma_{i}^{-1}J_{i}$ (see §4), which lowers the variance across all $N$ and reduces reliance on strong regularization, thereby also trimming bias. In Fig. 7 the curves shift downward nearly in parallel as $R$ grows, and the minimum RMSE occurs in the same $N$ band: discretization bias is governed by element size, so added sensors chiefly compress the variance without moving the bias-controlled optimum far to the right.
In Fig. 7, $\lambda$ is selected once per $R$ by the Quasi-Optimality criterion and then reused along the NN–sweep. This mimics practice (a single tuning per instrumentation layout). If $\lambda$ were re-tuned at every $(N,R)$, the right-hand rise would be partially damped (more regularization at larger $N$), but the qualitative U-shape set by (23) would remain.
The mid-span error is governed by the balance in (23): refine until discretization bias no longer dominates; beyond that point, without additional information (larger $R$ or richer load sets), the variance necessarily grows and RMSE worsens. This provides a practical criterion for choosing $N$ given a sensing plan.
@@ 6 Field application: openLAB research bridge
Fig.˜12 shows the reconstruction using six piecewise–constant elements. Each $EI$ value therefore represents an element average, and $75\%$ and $95\%$ credible bands quantify uncertainty. The posterior mean indicates lower flexural rigidity in span 1 than in span 2, consistent with construction records. While both spans possess identical cross-sections, Span 1 was cast with concrete strength class C25/30 and Span 2 with C50/60, implying a lower elastic modulus for the former. Crucially, the uncertainty is spatially heterogeneous: the bands remain narrow in the span interiors but widen significantly near the zero-moment zones, reflecting reduced sensor informativeness in those regions.